During the basketball season,Pierre scored 25% more points than Jim. Pierre scored 220 points for the season. The number of points jim scored was
Answer by funmath(2933)

During the basketball season,Pierre scored 25% more points than Jim. Pierre scored 220 points for the season. The number of points jim scored was
Let the amount of points that Jim scored be: x
Then 25% of x is: (25/100)x=.25x
"More than" means add:
x+.25x=220
1.25x=220
1.25x/1.25=220/1.25
x=176
:
Happy Calculating!!!